6/x+4/y=4
9/x-2/y=2

Answer by mananth(16946)

6/x+4/y=4
9/x-2/y=2

let 1/x=a
and 1/y =b
6a+4b=4.................(1)
9a-2b=2.................(2)

multiply (20 by 2
we get 18a-4b=4...........(3)
add (1) & (3)
24a=6
a= 6/24
a= 1/4
x=4
plug a in (1)
6*1/4+4b=4
3/2 + 4b=4
4b=4-3/2
4b=5/2
b= 5/8
y= 8/5